OpenStudy (anonymous):

Bill is able to save $35/week after working part-time and paying his expenses. These two formulas show his weekly savings: A. f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 B. f(n) = 35n 1. Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations. 2. Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get with this one?

OpenStudy (anonymous):

im still on part a

jimthompson5910 (jim_thompson5910):

So A. f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 you mean?

OpenStudy (anonymous):

yes, what do i do to get b?

jimthompson5910 (jim_thompson5910):

f(1) = 35 what's f(2) equal to?

OpenStudy (anonymous):

f(2)=70?

jimthompson5910 (jim_thompson5910):

f(n) = f(1) + f(n-1) f(2) = f(1) + f(2-1) f(2) = f(1) + f(1) f(2) = 35+35 f(2) = 70 correct

jimthompson5910 (jim_thompson5910):

f(3) = ???

OpenStudy (anonymous):

175?

jimthompson5910 (jim_thompson5910):

f(n) = f(1) + f(n-1) f(3) = f(1) + f(3-1) f(3) = f(1) + f(2) f(3) = 35+70 f(3) = 105 Hopefully you are seeing how recursion works?

OpenStudy (anonymous):

lol i dont know how i got 175

jimthompson5910 (jim_thompson5910):

maybe you did 105+70

jimthompson5910 (jim_thompson5910):

what is f(4)?

OpenStudy (anonymous):

175

jimthompson5910 (jim_thompson5910):

f(n) = f(1) + f(n-1) f(4) = f(1) + f(4-1) f(4) = f(1) + f(3) f(4) = 35 + 105 f(4) = 140

jimthompson5910 (jim_thompson5910):

always go back to the given recursive definition

OpenStudy (anonymous):

sorry i keep putting 70.

OpenStudy (anonymous):

are we on part 1 yet?

jimthompson5910 (jim_thompson5910):

1a, yes

jimthompson5910 (jim_thompson5910):

we have one more, f(5)

OpenStudy (anonymous):

175.00 finally

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

make sure to put that all in a table for 1a

OpenStudy (anonymous):

why am i making a table

jimthompson5910 (jim_thompson5910):

It says so 1. Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations.

OpenStudy (anonymous):

i didnt see that but when were done with 1&2 then we got to do the second half, i forgot to post it. :/

jimthompson5910 (jim_thompson5910):

well one thing at a time

jimthompson5910 (jim_thompson5910):

I guess if it's part of the problem (part of #1 or #2), then it needs to be included

OpenStudy (anonymous):

Income per week; Total after that week. $35.00 Week one: $35.00 $35.00 Week two: $70.00 $35.00 Week three: $105.00 $35.00 Week Four: $140.00 $35.00 Week Five: $175.00

OpenStudy (anonymous):

thats my table.?

jimthompson5910 (jim_thompson5910):

That looks correct

OpenStudy (anonymous):

III. calculations. Explain what your table means. IV. Use any formula of your choice to find f(40). Explain why you chose that method and what your answer means. Show your calculations. V. Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function.

jimthompson5910 (jim_thompson5910):

you could calculate f(40) the recursive way however that's very slow because you need f(1), f(2), f(3), ... etc etc all the way up to f(39) before you can even think of f(40)

jimthompson5910 (jim_thompson5910):

the quicker way is to use f(n) = 35n

jimthompson5910 (jim_thompson5910):

side note: notice how Week three: $105.00 and how f(n) = 35n f(3) = 35*3 f(3) = 105 So the function in part B is a much faster way to calculate the table and any large value of n

OpenStudy (anonymous):

So what number should i use for the up to 40 one?

jimthompson5910 (jim_thompson5910):

you'd use n = 40, but I recommend you use f(n) = 35n

jimthompson5910 (jim_thompson5910):

the method in part A is very slow to find f(40) or any large n value

jimthompson5910 (jim_thompson5910):

I'll brb

OpenStudy (anonymous):

so i times 35 by 40 everytime?

jimthompson5910 (jim_thompson5910):

no that's just to calculate f(40)

OpenStudy (driftracer305):

is that bill gates.......??